PDFTEX(1) PDFTEX(1) NAME pdftex,pdfinitex,pdfvirtex−PDFoutputfromTeX SYNOPSIS pdftex[options][commands] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. pdfTEXisaversionofTEXthatcancreatePDFfilesaswellasDVIfiles. ThetypicaluseofpdfTEXiswithapregeneratedformatsforwhichPDFoutputhasbeen enabled.ThepdftexcommandusestheequivalentoftheplainTEXformat,andthepdflatex commandusestheequivalentoftheLATEXformat. ThepdfinitexandpdfvirtexcommandsarepdfTEX’sanaloguestotheinitexandvirtexcommands.Inthisinstallation,theyaresymlinkstothepdftexexecutable. pdfTEX’shandlingofitscommand-lineargumentsissimilartothatofTEX. pdfTEXisbetasoftware. OPTIONS ThisversionofpdfTEXunderstandsthefollowingcommandlineoptions. --fmtformat Useformatasthenameoftheformattobeused,insteadofthenamebywhichpdfTEX wascalledora%&line. --helpPrinthelpmessageandexit. --iniBepdfinitex,fordumpingformats;thisisimplicitlytrueiftheprogramiscalledas pdfinitex. --interactionmode Setstheinteractionmode.Themodecanbeoneofbatchmode,nonstopmode,scrollmode, anderrorstopmode.Themeaningofthesemodesisthesameasthatofthecorresponding \commands. --ipcSendDVIorPDFoutputtoasocketaswellastheusualoutputfile.Whetherthis optionisavailableisthechoiceoftheinstaller. --ipc-start As--ipc,andstartstheserverattheotherendaswell.Whetherthisoptionisavailable isthechoiceoftheinstaller. --kpathsea-debugbitmask Setspathsearchingdebuggingflagsaccordingtothebitmask.SeetheKpathseamanual fordetails. --maketexfmt Enablemktexfmt,wherefmtmustbeoneoftexortfm. --no-maketexfmt Disablemktexfmt,wherefmtmustbeoneoftexortfm. --output-commentstring UsestringfortheDVIfilecommentinsteadofthedate. --prognamename Pretendtobeprogramname.Thisaffectsboththeformatusedandthesearchpaths. --shell-escape Enablethe\write18{command}construct.ThecommandcanbeanyBourneshellcommand.Thisconstructisnormallydisallowedforsecurityreasons. Web2C7.32x 29March1999 1 PDFTEX(1) PDFTEX(1) --translate-filetcxname Usethetcxnametranslationtable. --version Printversioninformationandexit. ENVIRONMENT SeetheKpathsearchlibrarydocumentation(the‘Pathspecifications’node)forprecisedetailsof howtheenvironmentvariablesareused.Thekpsewhichutilitycanbeusedtoquerythevalues ofthevariables. Onecaveat:InmostpdfTEXformats,youcannotuse˜ inafilenameyougivedirectlytopdfTEX, because˜ isanactivecharacter,andhenceisexpanded,nottakenaspartofthefilename.Other programs,suchasMETAFONT,donothavethisproblem. TEXMFOUTPUT Normally,pdfTEXputsitsoutputfilesinthecurrentdirectory.Ifanyoutputfilecannot beopenedthere,ittriestoopenitinthedirectoryspecifiedintheenvironmentvariable TEXMFOUTPUT.Thereisnodefaultvalueforthatvariable.Forexample,ifyousay texpaperandthecurrentdirectoryisnotwritable,ifTEXMFOUTPUThasthevalue /tmp,pdfTEXattemptstocreate/tmp/paper.log(and/tmp/paper.pdf,ifanyoutputis produced.) TEXINPUTS Searchpathfor\inputand\openinfiles.Thisshouldprobablystartwith‘‘.’’,sothat userfilesarefoundbeforesystemfiles.Anemptypathcomponentwillbereplacedwith thepathsdefinedinthetexmf.cnffile.Forexample,setTEXINPUTSto ".:/home/usr/tex:"toprependthecurrentdirecoryand‘‘/home/user/tex’’tothestandardsearchpath. TEXFONTS Searchpathforfontmetric(.tfm)files. TEXFORMATS Searchpathforformatfiles. TEXPOOL searchpathforpdfinitexinternalstrings. TEXEDIT Commandtemplateforswitchingtoeditor.Thedefault,usuallyvi,issetwhenpdfTEX iscompiled. FILES Thelocationofthefilesmentionedbelowvariesfromsystemtosystem.Usethekpsewhichutilitytofindtheirlocations. pdftex.pool EncodedtextofpdfTEX’smessages. texfonts.map Filenamemappingdefinitions. *.tfmMetricfilesforpdfTEX’sfonts. *.fmtPredigestedpdfTEXformat(.fmt)files. BUGS ThisversionofpdfTEXfailstotraparithmeticoverflowwhendimensionsareaddedorsubtracted. Caseswherethisoccursarerare,butwhenitdoesthegeneratedDVIfilewillbeinvalid. WhetherageneratedPDFfilewouldbeusableisunknown. pdfTEXisbetasoftware.Subscribetothepdftexmailinglistpdftex@tug.orgifyouintendto useit.Thisisamajordomolist,tosubscribesendamessagecontainingsubscribepdftexto Web2C7.32x 29March1999 2 PDFTEX(1) PDFTEX(1) majordomo@tug.org. SEEALSO tex(1),mf(1), AUTHORS TheprimaryauthorsofpdfTEXareHanTheThanh,PetrSojka,andJiriZlatuska. Web2C7.32x 29March1999 3